Solution x is 2% alcohol and solution y is 6% alcohol. A drugstore owner wants to mix them in order to get a 60 liter of solution that is 3.2% alcohol. How many liters of each should the owner use?
My dear tutor i would be very glad if you can help me solved this problem. I had been trying my very best but really iam poor in math. THANKS so much and more power........

Answer by longjonsilver(2297)

DEFINITION:
Let x = amount of solution needed.
Leap of common sense... how much of solution y is therefore needed? Answer is "the rest of the 60litres", so we have (60-x).

BEFORE:
0.02*x + 0.06(60-x)

AFTER:
0.32*60

and the before must equal the after surely? :-) so...
0.02*x + 0.06(60-x) = 0.032*60
0.02x + 3.6-0.06x = 1.92
-0.04x = -1.68
x = (-1.68)/(-0.04)
x = 42 litres.

And so 60-42 is 18.

x=42litres
y=18 litres

jon.